Experiments

Last modified by Michael Sibayan on 2015/02/13 12:32
Comments (0) · Attachments (0) · History

Category: Experiments (10 posts) [RSS]

Jul 29 2020

Gravity flow experiment

Gravity flow experiment, July 2020

Prior to the experiment the NovaLynx tipping buckets were calibrated (see this post).

The experiment was repeated separately for each hillslope, since the storage tanks cannot hold enough water for a 12h rain with the Multi 1 circuit on all 3 hillslopes simultaneously.

For each hillslope a 1h rain event was performed to calibrate the rainfall distribution followed by a 12h rain event on the following day to reach a gravity flow regime down to a depth of 50cm.

Details on the rainfall calibration: rainfall_calibration.xlsx

Schedule: 

Date Task Comment 
 CENTER 
 2020-07-06 Spanned ropes on Center hillslope to mark cup locations for rainfall calibration. Masts were raised in the morning. Kept raised until after the 12h rain event.
 Ropes spanned upslope-downslope with pre-labeled marks every 2m from 0-30m.
 2020-07-07 Placed cups on Center hillslope and removed ropes. 
 2020-07-08 Issues with Tipping Bucket counts, switched to cumulative counts to reduce impact. See also NovaLynx calibration
 2020-07-09 Rainfall calibration, Multi 1, Center hillslope.
 Collected and weighed cups after 1h of rain (08:30-09:30).
 Details in attached excel file. Air handler off during the rain.
 2020-07-10 12h rain event (8:30-20:30), Multi 1, Center hillslope.  Air handlers were turned off during rain.
Discharge collection for Center -1 (PE102 and NovaLynx) was leaking. Fixed at 19:40.
 Mast lowered after the rain event.
 EAST 
 2020-07-13 Spanned ropes on East hillslope to mark cup locations for rainfall calibration. Masts were raised in the morning. Kept raised until after the 12h rain event.  
 Ropes spanned upslope-downslope with pre-labeled marks every 2m from 0-30m.
 2020-07-14 Placed cups on East hillslope and removed ropes. 
 2020-07-15 Rainfall calibration, Multi 1, East hillslope.
 Collected and weighed cups after 1h of rain (08:30-09:30).
 Details in attached excel file. Air handler off during the rain.
 2020-07-16 12h rain event (8:30-20:30), Multi 1, East hillslope. Air handlers were turned off during rain.
 Mast lowered after the rain event. 
 WEST 
 2020-07-20 Spanned ropes on West hillslope to mark cup locations for rainfall calibration. Masts were raised in the morning. Kept raised until after the 12h rain event.  
 Ropes spanned upslope-downslope with pre-labeled marks every 2m from 0-30m.
 2020-07-21 Placed cups on West hillslope and removed ropes. 
 2020-07-22 Rainfall calibration, Multi 1, West hillslope.
 Collected and weighed cups after 1h of rain (08:30-09:30).
 Details in attached excel file. Air handler off during the rain.
 Gantry malfunction caused an interruption of several hours during the cup collection.
 2020-07-23 12h rain event (8:30-20:30), Multi 1, West hillslope. Air handlers were turned off during rain.
 Mast lowered after the rain event. 

After the experiment, cups, labeled ropes and cardboard transport boxes used for the cup collection during the rainfall calibration were stored in upper habitat office 10.

Aug 01 2017

LEO_W_broken_sprinklerhead

Broken sprinkler head on multi 2 upon opening valve 8/1/17 7:00. Switched to Multi 1 circuit and will rain with Multi 1 today

Jan 18 2017

bigPERTH_SoilCoring_End of Experiment

Second coring conducted on 12/21/2016.

Modifications included:

1. Battery-operated drill head (coring happened faster because of this)

2. Coring locations were 1 (26.75,0), 2 (17.75, 0), 3 (9.75, 0), 4 (3.75, 0), 5 (9.75, 3), 6 (9.75, -3)

3. Cores recovered were shorter in length as compared to first coring event on November 8th. The depth of the soil core in the hillslope was however 85 cm. So either the soil is compacted or getting displaced. 

4. All cores were divided into 5 equal sub-parts. 

5. Katerina has information on volumteric dry weight for this set of cores.

Nov 15 2016

PE102-075 flow meter added to Center Bay Leak collector

Added a Seametrics PE102-075 flow meter inline with the Leak tipping bucket in the Center Bay. 

Valid flow rates for this model are 0.2 - 20 GPM. 

Data for Leak tipping bucket between 11/15/2016 11:40 - 13:06 are not valid

Nov 14 2016

Seepage clogged

Seepage clogged and fixed,flow rates down while fixing ...

Nov 11 2016

bigPERTH_SoilCoring_Beginning of Experiment

Soil cores were collected from all three hillslopes on 11/8/2016.

  • See attached document for core location and nomenclature
  • 6 cores per slope * 3 slopes = 18 cores

    **Once coring is done, fill gaps with blue barrel soil material that have received the same volume of total water as the big hillslope.SoilCore_bigPERTH.jpgSoilCore_bigPERTH.jpg

    Coring (Nate, Aaron); LEO spaceProcessing (Ed/Aditi); Analytical LabStoring/Archiving (Ed/Aditi)

    1.    Core at 6 locations per slope

    2.    Tape core ends shut with electrical tape, 

    label cores with labeling tape to indicate surface and deepest end. 

    Store cores in cooler prior to processing

    3.    Hand over cores to Ed/Aditi

    4.     Measure cores for sub-sections

    5.     Cut cores in the lab

    6.     Place in labeled bags, homogenize

    7.     Aliquot for DNA, RNA in microcentrifuge tubes (plus archive in 50 ml tubes),

    8. Flash freeze microbio samples, drop in dry ice till stored at -80

    9. Ed takes care of geochem samples

  • Update on sampling:
  • Central slope was cored first,starting at approximately 10.45 AM-12.00PM. Followed by west slope (starting at 1.30 PM to 3 PM), followed by east slope (3.45-5.30 pm). Central slope cores were collected one at a time (Cores were stored on dry ice while being collected). West and East slope cores were collected as one set (Nate and Aaron did not hand down each core prior to collecting the next one. So east and west cores were not stored in cooler while being collected). The metal core has to be cleaned every time with water so for east and west cores, wash bottle was carried on the gantry along with a bucket to wash the cores. Also, location 2-6 are cored at one go, for location 1, power drill connection has to be pulled from the top of the slope. 
  • For next coring, buy big cooler, add ice, drop cores in it. Ask Nate and Aaron to carry cooler with them on gantry. 
  • The average core length was 65-70 cm. Four 15 cm sub-depths were collected from each core plus a fifth depth sample that comprised of whatever was left over.
  • Sub-samples were homogenized in whirl pack bags, Ed took C and N samples from the bags, bag handed to Aditi,  who subsampled into centrifuge tubes. Started out with sampling into 1.5 ml centrifuge tubes. Stopped after central slope. Way too much time was being spent in doing this (the integrity of the cores was in question since they were out in bags now). Rest of the samples are in whirlpack bags and 50 ml centrifuge tubes. All bags and tubes were flash frozen and stored on dry ice, then brought back to campus and stored at -80 C at Raina Maier's lab.  An alternate is to have more people work to make sure this gets done faster. Plus have another cooler of ice to store homogenized soil bags in before sub-sampling.

Nov 07 2016

LEO bigPERTH experiment started

The LEO bigPERTH experiment was started on Nov. 6, 2016 at 23:00 MST.

Irrigation actuators on LEO East and West did not respond, so circuits were opened and closed in manual mode with a slight start-time delay (1-5 minutes).

Jul 24 2016

LEO Sampling Interval Change

July 24, 2016
The sampling interval for most LEO sensors, excluding 5TM, MPS-2, and CS451, has been set to 1 minute for a series of above slope air flow experiments. Since there is no anticipated irrigation I have decided to leave the SDI-12 sensors (listed above) at the standard 15 minute logging interval.

Feb 19 2013

LEO-E steady state experiment 1

At approximately 10:00 MST on February 18, 2013, irrigation circuit Multi2 was activated to initiate the steady state rain experiment.
At approximately 7:45 MST on February 19, 2013, irrigation was stopped.

Jan 02 2013

LEO-E rain drop analysis

On 1/2/13 between approx 10 am and 12:30 pm, rain drop diameter and velocity distributions were measured on LEO1 slope. Measurements were acquired at two locations (upper and lower slope along central axis) for each rain circuit and then for all circuits simultaneously for 5 minutes at each setting.
(Notes from Jon Pelletier)

Tags:
Created by Michael Sibayan on 2015/02/13 12:32

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 6.0.1 - Documentation